OpenStudy (anonymous):

PLEASE HELP. TEST TOMORROW. Look at the recursive formula. A1=6 for n>1,An=A(n-1)+4 What are the first 4 terms of this sequence? A. 5,6,7,8 B. 6,10,14,18 C. 9,12,15,18 D. 10,14,18,22 how do i solve this??

OpenStudy (anonymous):

you have \(a_1=6\) right? so next step is to compute \(a_2\)

OpenStudy (anonymous):

\[a_2=a_1+4=6+4=10\] for the next term

OpenStudy (anonymous):

one after that is \[a_3=a_2+4=10+4=14\] etc etc

OpenStudy (anonymous):

is that all i have to do?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you